Your classes are being scheduled for the upcoming school year. You will be taking English, Math, History, PE, Spanish, Woodshop and Computer. What is the probability that Workshop will be scheduled during 1st period and Math during 2nd period?
Answer by stanbon(75887)

Your classes are being scheduled for the upcoming school year. You will be taking English, Math, History, PE, Spanish, Woodshop and Computer. What is the probability that Workshop will be scheduled during 1st period and Math during 2nd period?
----
# of ways to succeed:
1st period : 1 way
2nd period : 1 way
3rd thru 7th: 5! ways
==========================
# of random arrangements: 7!
---
P(order as discribed) = (1*1*5!)/7! = 1/(7*6) = 1/42
================
Cheers,
Stan H.
==========================